concert tickets: tickets for a concert cost $5, $3, and $2. twice as many $5 tickets were sold as $2 tickets. the recipts for 750 tickets were $2625. haw many of each price ticket were sold? [i have no clue how ti set this up]
Answer by scott8148(6628)

let x="$2 tickets", so 2x="$5 tickets" and 750-3x="$3 tickets"

2(x)+5(2x)+3(750-3x)=2625 __ 2x+10x+2250-9x=2625 __ 3x=375